### v1

# Why Reasoning Matters? A Survey of Advancements in Multimodal Reasoning (v1)

Abstract
Reasoning is central to human intelligence, enabling structured problem-solving across diverse tasks. Recent advances in large language models (LLMs) have greatly enhanced their reasoning abilities in arithmetic, commonsense, and symbolic domains. However, effectively extending these capabilities into multimodal contexts—where models must integrate both visual and textual inputs—continues to be a significant challenge. Multimodal reasoning introduces complexities, such as handling conflicting information across modalities, which require models to adopt advanced interpretative strategies. Addressing these challenges involves not only sophisticated algorithms but also robust methodologies for evaluating reasoning accuracy and coherence. This paper offers a concise yet insightful overview of reasoning techniques in both textual and multimodal LLMs. Through a thorough and up-to-date comparison, we clearly formulate core reasoning challenges and opportunities, highlighting practical methods for post-training optimization and test-time inference. Our work provides valuable insights and guidance, bridging theoretical frameworks and practical implementations, and sets clear directions for future research.
1 Introduction
Reasoning is a fundamental aspect of human intelligence, enabling us to solve complex problems effectively. In LLMs, reasoning has shown promising capabilities, leading to significant advancements in various domains (Bi et al., 2024b), including arithmetic, commonsense, and symbolic reasoning. For instance, Chain-of-Thought (CoT) prompting mimics human stepwise problem-solving to boost LLM performance, especially in very large models.
Beyond CoT, other prompting strategies, such as generated knowledge prompting and tree search algorithms (e.g., STAR search (Zelikman et al., 2022)), have been incorporated into LLM training to further enhance reasoning capabilities. These methods encourage the model to articulate its internal reasoning in natural language, which reinforces its understanding of the task. By laying out intermediate steps, the model can verify and adjust its logic before arriving at the final answer.
Figure 1: Papers on visual reasoning per quarter over the last three years, with state computed using referenced papers. (Data current to mid-March 2025.)
While reasoning in LLMs has progressed significantly, extending these capabilities to multimodal tasks remains an emerging frontier, as multimodal large language models (MLLMs) must navigate the added complexity of interpreting and integrating information from both visual and textual modalities, often resolving ambiguities, inconsistencies, or gaps that arise when the two sources conflict or diverge.
Effective reasoning is central to enabling MLLMs to achieve a compositional understanding—allowing them to deconstruct complex tasks into interpretable, modality-spanning steps. It also supports iterative processes such as error correction and self-refinement, where models can revise their outputs by reevaluating both visual and textual cues.
Moreover, strong reasoning capabilities help clear up confusion by inferring spatial relationships, handling counterfactuals, and choosing appropriate tools or actions. These mechanisms also play a crucial role in mitigating hallucinations by grounding model outputs in cross-modal evidence, ultimately improving both accuracy and trustworthiness. Finally, reasoning extends the potential of MLLMs to handle hypothetical scenarios, helping them look ahead and explore “what-if” situations that span both language and vision.
Initial research has demonstrated that integrating Chain-of-Thought (CoT) reasoning across different modalities can significantly enhance the performance of MLLMs. For example, Yao et al. (2024); Zhang et al. (2024a); Bi et al. (2024a) reports that the LLama series models’ reasoning abilities are greatly improved by solely utilizing filtered self-generated reasoning paths.
Building on this momentum, our work provides a comprehensive overview of reasoning in both LLMs and MLLMs, with a focus on its techniques, applications, and future directions. We aim to bridge the gap between text-based and multimodal reasoning, offering insights into how reasoning can further enhance the capabilities of LLMs in multimodal contexts. In the following sections, we begin with a clear problem formulation and definition of reasoning, followed by an in-depth analysis of reasoning techniques—from post-training strategies to test-time computation. We then examine recent trends in datasets and benchmarking. Our goal is to offer a well-organized and accessible survey that supports both theoretical understanding and practical implementation of reasoning in LLMs and MLLMs.
2 Background
In complex question-answering tasks, directly predicting an answer can be highly uncertain due to the vast range of possible responses. A more effective approach involves breaking down the reasoning process into a sequence of intermediate steps. This structured method not only improves interpretability but also helps reduce uncertainty at each inference step.
Formally, let $Q$ denote a given question. Conventional language models typically aim to model the conditional probability $P(A\mid Q)$ of the answer $A$ given the question. However, as the complexity of $Q$ increases, the prediction becomes more uncertain—reflected by a rise in the conditional entropy $H(A\mid Q)$. A natural approach to mitigate this uncertainty is to decompose the reasoning process into a sequence of intermediate steps. This leads to modeling the answer generation as a structured chain of conditional probabilities:
$P(\text{Step}_{1}\mid Q)\cdot P(\text{Step}_{2}\mid\text{Step}_{1},Q)\cdot\dots\cdot P(\text{Step}_{t}\mid\text{Step}_{1:t-1},Q)\cdot P(A\mid\text{Step}_{1:t},Q)$ (1)
Such a decomposition encourages the model to reduce uncertainty at each stage and improves interpretability and robustness in complex question answering. This step decomposition is useful because conditioning on prior steps reduces uncertainty. Specifically, the conditional entropy satisfies: $H(\text{Step}_{t}\mid\text{Step}_{<t},Q)\leq H(\text{Step}_{t}\mid Q)$, meaning each step becomes easier to predict as more context accumulates. From a probabilistic standpoint, the entire reasoning process can be viewed as a trajectory, denoted as $\tau=(\text{Step}_{1},\dots,\text{Step}_{t})$, which represents a complete reasoning path. Using this, the overall probability of an answer given a question can be written as a marginal over all possible trajectories:
$P(A\mid Q)=\sum_{\tau}P(A\mid\tau,Q)\cdot P(\tau\mid Q).$ (2)
However, this formulation presents a challenge: enumerating all possible reasoning paths $\tau$ is intractable. In practice, CoT reasoning offers a practical workaround by sampling a single trajectory $\hat{\tau}$ from the distribution $P(\tau\mid Q)$, and using it to approximate the answer:
$P(A\mid Q)\approx P(A\mid\hat{\tau},Q),\quad\text{where}\quad\hat{\tau}\sim P(\tau\mid Q).$ (3)
This is beneficial for many tasks if the model is capable of generating one coherent reasoning path that captures the correct structure of the solution, thereby reducing uncertainty and improving answer accuracy. More advanced methods such as Tree-of-Thought (ToT) expand on CoT by generating and evaluating multiple reasoning paths, and can be seen as a multiple point estimation approach. In this case, the model generates a set of reasoning paths $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ and evaluates them to select the most promising one.
Given the challenge of enumerating all possible reasoning paths, our objective becomes clear: to generate and identify only the most promising trajectories. Achieving this effectively requires two key components: (i) Generation – a strong model capable of producing high-quality reasoning paths $\tau$; (ii) Selection – a reliable mechanism for evaluating and choosing the best candidate path.
These components are deeply interconnected: improved generation policies yield better trajectories, which in turn provide stronger supervision signals to further refine the model. This creates a virtuous cycle of mutual enhancement, as illustrated in Figure 2.
Figure 2: Framework illustrating training and inference for reasoning optimization. A virtuous cycle emerges as better policies generate improved trajectories, which in turn enhance the model through stronger supervision.
At inference time, we want to search for the best $\tau$ under the current model. During training, we aim to improve the model so that it generates reasoning paths that are more likely to lead to correct answers.
Broadly speaking, recent approaches fall into two main categories:
Post-training improvement involves optimizing the model policy $\pi$ to better align outputs with desired utility, typically through techniques like fine-tuning or reinforcement learning. Formally, this corresponds to solving Equation (4), where the goal is to learn a policy that maximizes expected utility over generated outputs $\tau$.
Test-time compute focuses on improving LLM performance during inference without modifying the model’s core parameters. This aligns with the objective in Equation (4), where the goal is to select the most useful output trajectory $\tau$ under a fixed policy $\pi^{*}$, optimizing accuracy, coherence, and efficiency through dynamic strategies like CoT, ToT, and search-based reasoning.
$\begin{cases}\pi^{*}=\displaystyle\arg\max_{\pi}\,\mathbb{E}_{\tau\sim\pi}[\text{Utility}(\tau)]&\text{(Training: Learn a better model)}\\ \tau^{*}=\displaystyle\arg\max_{\tau\sim\pi^{*}}\,\text{Utility}(\tau)&\text{(Inference: Select best trajectory)}\end{cases}$ (4)
In the following sections, we will discuss these two components in detail.
3 Post-training improvements
To enhance the quality of reasoning, we aim to learn a model that generates high-utility reasoning trajectories. One effective approach frames this as a Markov Decision Process (MDP), where the model is trained to maximize the expected return of a reasoning path $\tau$:
$\max_{\theta}\ \mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t}\gamma^{t}R(s_{t},a_{t})\right]\quad\text{with}\quad\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})A(s_{t},a_{t})\right]$ (5)
$s_{t}$ is the sequence of tokens up to time $t$, and $a_{t}$ is the next token. The advantage function $A(s_{t},a_{t})$ estimates the benefit of choosing $a_{t}$ in state $s_{t}$, guiding the model toward higher-utility token sequences. This objective can be directly optimized with respect to the model parameters using policy gradient methods (Williams, 1992). Notably, this formulation enables the training of LLMs not merely to predict the next token, but to generate complete reasoning trajectories that are optimized for long-term rewards.
[…]
3.1 Policy Optimization
To solve the above reward maximization problem, recent advancements in visual reasoning have prominently focused on integrating reinforcement learning (RL) and imitation learning (IL) to align MLLMs more closely with human reasoning. IL approaches, such as Thought Cloning (Wei et al., 2025), surpass traditional cloning by aligning intermediate reasoning steps instead of final outputs alone. Coupled with iterative methods like DAgger, this reduces dataset shift and hallucination. LLaVA-Critic (Xiong et al., 2024) further generalizes reward signals via preference-based alignment, boosting multimodal reasoning effectiveness with minimal modality-specific fine-tuning. RL from Simulations (RLS3) (Waite et al., 2025), employing Soft Actor-Critic (SAC) agents, significantly enhances spatial reasoning accuracy and efficiency in models like PaliGemma, outperforming random exploration baselines (Waite et al., 2025). FuRL (Fu et al., 2024c) addresses RL reward misalignment through dual-stage alignment, effectively mitigating sparse reward challenges. Adaptive Reward Programming (ARP-DT+) (Kim et al., 2023), integrating CLIP representations with Decision Transformers, enhances reasoning robustness across domains by employing active inference to counter distractions and domain shifts.
3.2 Reward Alignment
Recent approaches have explored iterative refinement and reflection mechanisms, token-level rewards, and automated benchmarking to enhance reasoning quality and mitigate hallucinations. PARM++ (Guo et al., 2025) introduces a reflection-based mechanism that iteratively refines outputs by identifying misalignments and actively correcting them through self-assessment loops, significantly improving test-time GenEval scores up to three reflection cycles (Guo et al., 2025). Similarly, REVERIE employs rationale-based training, explicitly supervising reasoning steps, thus improving coherence and reducing hallucinations compared to models without rationales (Zhang et al., 2024f). Fine-grained reward models, such as MMViG (Yan et al., 2024) and Reward Alignment via Preference Learning (RAPL) (Tian et al., 2024), emphasize granular feedback at individual reasoning steps, achieving precise improvements without manual reward tuning. Similarly, CLIP-DPO (Ouali et al., 2024) leverages pre-trained CLIP directly as a reward function, streamlining reward modeling. Challenges such as constrained perceptual fields are addressed through dynamic evaluation strategies like zooming and shifting, emphasizing the importance of sequential decision-making in visual reasoning (Wang et al., 2024j). Structural issues like positional bias and patch-boundary effects further underscore the need for targeted architectural optimizations (Zhang et al., 2024d). Recent approaches integrate reward signals directly into generation. TLDR (Fu et al., 2024a) applies token-level binary rewards, enabling real-time feedback to mitigate hallucinations and boost both annotation efficiency and backbone model performance (Fu et al., 2024a). Complementarily, EACO enhances intermediate object recognition, significantly reducing hallucinations compared to earlier models like LLaVA-v1.6-7B (Wang et al., 2024g). Distillation frameworks also contribute: FIRE leverages iterative student–teacher feedback and structured evaluation to refine visual reasoning (Li et al., 2024d), while SILKIE demonstrates the scalability of GPT-4V-generated feedback in improving perception and cognition without human intervention (Li et al., 2023b).
3.3 Model Architecture
Nested architectures, such as MaGNeTS, have emerged as effective means of balancing computational efficiency with model accuracy by employing parameter-sharing and caching mechanisms during inference (Goyal et al., 2025). Intrinsic activation approaches, notably ROSS, bypass external modules by integrating multimodal understanding directly into the model’s core, yielding improved adaptability to diverse visual inputs, including depth maps (Wang et al., 2024c). Additionally, DIFFLMM incorporates diffusion models and attention-based segmentation, enhancing visual grounding precision (Cao et al., 2024a), while Mini-Monkey targets the ”sawtooth effect” in lightweight models through adaptive cropping and scale compression, achieving superior visual-text alignment without extensive computational resources (Huang et al., 2024b). SEA dynamically adjusts embedding alignment strategies to maintain performance across varying resolutions and model sizes (Yin et al., 2024). Concurrently, PAE-LaMM demonstrates the combined effectiveness of vision encoder adaptation and pixel reconstruction tasks, systematically enhancing visual detail awareness and question-answering performance (Gou et al., 2024). Conversely, LLAVIDAL and EventLens uniquely integrate domain-specific visual elements, such as object-centric embeddings and event-aware tokens, respectively, directly within LLM architectures to effectively reason about activities and temporal contexts (Reilly et al., 2024; Ma et al., 2024a).
3.4 Spatial Temporal Modeling
Unlike pure language reasoning, visual reasoning involves a richer action space, allowing operations such as zooming, cropping, and frame selection, which enables the model to interact with and manipulate visual inputs. The GeoGLIP pipeline utilizes geometric pre-training with a dynamic feature router to advance fine-grained visual mathematical understanding by balancing symbolic and visual modalities based on task demands (Zhang et al., 2025). Concurrently, hallucination issues are mitigated via targeted fine-tuning and prompting strategies grounded in recent vision-language work like CapQA (Hu et al., 2025). For spatial reasoning, simpler graph topologies have been shown to enhance performance, while structural complexity reduces accuracy in edge-centric tasks (Zhu et al., 2024). Augmentation via uncertainty-aware active inference further boosts attribute detection precision (Zhang et al., 2024e). Intrinsic spatial-temporal modeling is advanced through architectures embedding multi-modal understanding without reliance on external depth tools (Wang et al., 2024c), and dual-branch structures improve video temporal grounding in multi-hop reasoning (Chen et al., 2024f). The shift toward multimodal benchmarks better reflects real-world spatial-temporal reasoning (Li et al., 2024a). Integrating rationalization with answer generation enhances temporal reasoning (Zohar et al., 2024), while automated annotation improves data scalability and bridges gaps between commercial and open-source Video-LLMs (Li et al., 2024e). Multi-modal integration and active inference optimize modality selection for multi-step spatial-temporal reasoning (Zhang et al., 2024l). Adjusting learning objectives to minimize hallucinations improves precision and EOS decision-making (Yue et al., 2024).
Figure 3: Search framework where language models explore and refine reasoning paths. Trajectories are scored using reward models, based on expected utility or final output quality, and guided by feedback, world models, and evaluators to select the most promising steps.
4 Test-Time compute
At test time, the goal is to find the reasoning trajectory $\tau$ that maximizes a utility function as shown as red and green path in Figure 3
$\tau^{*}=\arg\max_{\tau}\,\mathcal{U}(\tau),\quad\text{where }\mathcal{U}(\tau)=\begin{cases}\sum_{t}r(s_{t},a_{t})&\text{(MDP-style reward)}\\ \mathbb{P}(\text{goal}\mid\tau)&\text{(Goal likelihood)}\\ \mathbb{P}(\tau\succ\tau^{\prime})&\text{(Preference-based)}\\ f(\text{rank}(\tau))&\text{(Ranking-based)}\\ \vdots&\text{(Others: risk-sensitive, etc.)}\end{cases}$ (6)
The utility $\mathcal{U}(\tau)$ can be defined in various ways, including cumulative rewards (MDP-style), goal likelihood, preference comparisons, and ranking-based scores. The choice of utility depends on the supervision available and the nature of the task. The core challenge remains sampling high-utility trajectories under the chosen formulation.
Since directly modeling $\mathcal{U}(\tau)$ is often difficult, we use reward models to approximate it. These models can include (1) absolute rewards (as in standard RL), (2) pairwise preference models (e.g., DPO), and (3) ranking-based models.
[…]
4.1 Search Strategies
Monte Carlo Tree Search (MCTS) has emerged as a powerful tool for managing uncertainty and sequential decision-making. Imam et al. (2025) applied MCTS to enhance temporal reasoning, reducing hallucinations in time-dependent data. Building on this, CoMCTS(Yao et al., 2024) unified multiple MLLMs within a collaborative tree search, addressing model bias and improving accuracy and efficiency. Vision-specific strategies advanced with VisVM (Wang et al., 2024f), which estimates long-term candidate value to reduce hallucinations, outperforming immediate-alignment methods like CLIP-PRM. Video reasoning similarly benefited from reward modeling and adaptive exploration to prioritize key frames (Yang et al., 2024b). Structured search methods also gained traction. LLaVA-o1 (Xu et al., 2025) uses stage-level beam search for complex reasoning, while MVP (Qu et al., 2024) aggregates certainty across multi-perspective captions to resist adversarial inputs. DC2 (Wang et al., 2024e) applies MCTS-based cropping to focus on salient image regions for high-resolution reasoning. Multimodal and temporal search frameworks like VideoVista (Li et al., 2024e), WorldRetriver (Zhang et al., 2024l), and DynRefer (Zhao et al., 2024) surpass static baselines using adaptive sampling, fusion, and stochastic inference. Step-by-step comparative reasoning with LMs also enhances video QA (Nagarajan & Torresani, 2024). Innovations in context refinement (VURF (Mahmood et al., 2024)), image decomposition (V*(Wu & Xie, 2023)), and dynamic tool use (AVIS(Hu et al., 2023)) highlight the shift toward adaptive visual reasoning. FuRL (Fu et al., 2024c) aligns reward modeling with iterative fine-tuning for better performance. In embodied AI, combining chain-of-thought, self-verification, and MCTS-driven planning enables scalable, robust decision-making in dynamic environments (Shin et al., 2024).
4.2 Adaptive Inference
Adaptive inference is reshaping vision-language reasoning by enabling dynamic, context-sensitive processing that improves both accuracy and efficiency. Central to this is the iterative evaluation and refinement of outputs, often through internal feedback and external verification. LOOKBACK enhances correction accuracy through iterative visual re-examination within each reasoning step (Wu et al., 2024b). IXC-2.5 balances performance and response length via Best-of-N sampling guided by reward models (Zang et al., 2025), while PARM++ uses reflection-based refinement to align generated images with prompts (Guo et al., 2025). Further innovations include ProgressCaptioner’s sliding-window approach for tracking action progress over time (Xue et al., 2024). To combat multi-modal hallucinations, MEMVR triggers visual retracing based on uncertainty, optimizing cost-accuracy tradeoffs (Zou et al., 2024a), while MVP aggregates certainty across diverse views (Qu et al., 2024). SID applies token-level contrastive decoding to filter irrelevant content (Huo et al., 2024). Models like DualFocus integrate macro and micro reasoning for fine-grained attention control (Cao et al., 2024b), and LISA++ enables test-time compute scaling without retraining (Yang et al., 2024a). PerceptionGPT accelerates inference via adaptive token embeddings that encode spatial cues and dynamically weigh layers (Pi et al., 2023).
4.3 Reward
Guo et al. (2025) introduce ORM and PRM, showing that ORM’s final-step evaluation significantly enhances image generation through ranking data. PARM++ further refines this method by enabling iterative reflection for self-correction, thus improving prompt fidelity. Complementarily, Hao et al. (2025) highlight persistent errors in VLM spatial reasoning and explanation, proposing reward re-ranking as an interim solution. To assess VL-GenRMs more comprehensively, Li et al. (2024b) propose VL-RewardBench, pinpointing perception and reasoning challenges while advocating critic training and co-evolutionary reward learning. Self-training methods also gain prominence: R3V (Cheng et al., 2024b) leverages synthesized rationales to enhance noise-robust Chain-of-Thought (CoT) refinement, while CECE (Cascante-Bonilla et al., 2024) enriches evaluation using LLM-generated entailment/contradiction captions. Enhanced supervision methods also emerge. Zhang et al. (2024k) enhance CoT reasoning through ShareGPT-4O distillation, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). LLaVA-Critic (Xiong et al., 2024) builds upon this using critic-guided DPO without external human feedback. FuRL (Fu et al., 2024c) introduces a two-stage tuning approach addressing misalignments and temporal errors, while HYDRA (Ke et al., 2024) employs dynamic loops for adaptive instruction ranking. Additionally, m&m’s (Ma et al., 2024b) provides a structured framework for evaluating multi-step planning via tool and argument accuracy.
4.4 Feedback
VOLCANO employs a three-stage iterative framework—initial response generation, visual self-assessment, and revision—to reduce hallucinations by emphasizing accurate image details (Lee et al., 2024). Similarly, LOOKBACK mandates explicit atomic verification against images to enhance both critique and correction processes for LVLMs (Wu et al., 2024b). LLaVA-Critic advances feedback by incorporating diverse critic instructions and iterative DPO training, leveraging internal data (Xiong et al., 2024). Additionally, tailored visual prompting with iterative binary verification effectively enhances semantic grounding in models like LLaVA-1.5, ViP-LLaVA, and CogVLM (Liao et al., 2024a). Minimalist reinforcement learning setups have shown effectiveness, particularly for multimodal mathematical reasoning, with larger models like InternVL2.5-38B benefiting from difficulty-based data filtering to stabilize training (Meng et al., 2025). Explicit visual CoT mechanisms, exemplified by VisCoT, integrate bounding box predictions and iterative cropping to enhance visual question-answering performance (Shao et al., 2024). In video understanding, VIDEOTREE utilizes adaptive hierarchical clustering for efficient navigation through structured tree-based representations (Wang et al., 2024i). Lastly, prompt engineering strategies like MiHO and MiHI significantly reduce hallucinations without model retraining, post-processing evaluations with GPT-4 further enhancing reliability without architectural modifications (Han et al., 2024; Hu et al., 2025).
4.5 Iterative Refinement
Adaptive inference transforms vision-language reasoning by facilitating dynamic, context-sensitive processing to enhance accuracy and efficiency. Central to this approach is iterative output refinement through internal feedback and external verification. For instance, IXC-2.5 employs Best-of-N sampling guided by reward models to optimize performance and response length (Zang et al., 2025), while PARM++ uses reflection-based refinement for better alignment between images and prompts (Guo et al., 2025). Further advancements include LLaVA-o1’s structured perception and stage-level beam search for decomposing complex reasoning tasks (Xu et al., 2025). To mitigate multi-modal hallucinations, MEMVR implements visual retracing based on uncertainty to optimize the cost-accuracy tradeoff (Zou et al., 2024a), MVP aggregates certainty across diverse views (Qu et al., 2024), and SID employs token-level contrastive decoding to eliminate irrelevant content (Huo et al., 2024). Moreover, DualFocus integrates macro- and micro-level reasoning for precise attention control (Cao et al., 2024b), LISA++ enables test-time compute scaling without retraining (Yang et al., 2024a), and PerceptionGPT accelerates inference through adaptive token embeddings encoding spatial cues and dynamically weighted layers (Pi et al., 2023).
4.6 Dataset Curation
Recent work shows that high-quality, strategically curated datasets outperform large but noisy alternatives in guiding reasoning paths. R-CoT (Deng et al., 2024) enhances geometric reasoning via reverse generation and stepwise synthesis. Task-specific datasets such as SMIR (Li et al., 2025a) show up to 8% performance gains over generic datasets, demonstrating the value of targeted curation. Gradual complexity in dataset construction, as shown in VIREO (Cheng et al., 2024a), ensures foundational reasoning skills before introducing advanced tasks. Scalable dataset creation through automation also proves effective. VideoVista (Li et al., 2024e) uses auto-annotations for video reasoning, outperforming manually curated sets. DecoVQA+ (Zhang et al., 2024b) explicitly teaches when to apply question decomposition, aided by SelectiveVQD loss for better strategy selection. LocVLM (Ranasinghe et al., 2024) scales with pseudo-data and implicit feedback signals like location and relevance prediction. CogCoM (Qi et al., 2024) incorporates tasks like grounding and manipulation to reinforce step-by-step reasoning.
5 Datasets and Benchmarks
Structured and Task-Specific Reasoning Datasets such as Visual-RFT(Liu et al., 2025c), CapQA(Hu et al., 2025), GUIDE(Liang et al., 2024a), STAR(Wu et al., 2024a), Visual Genome(Zhang et al., 2024f), VL-GPT(Zhu et al., 2023), and Interfacing (Zou et al., 2024b) have introduced structured prompting that significantly enhances models’ reasoning accuracy by explicitly guiding reasoning processes. In contrast to earlier benchmarks like STAR, these datasets place a stronger emphasis on task-specific reward functions and structured inference, aiming to improve generalization across diverse visual and linguistic contexts.
Temporal and Spatial Reasoning Temporal benchmarks like VisualQA, and TemporalVQA (Imam et al., 2025), TLQA (Swetha et al., 2025), REXTIME (Chen et al., 2024c), FrameCap (Xue et al., 2024), and VideoVista (Li et al., 2024e) have revealed substantial limitations in current multimodal language models, achieving accuracy significantly below human levels (15% versus 90%). These benchmarks emphasize the necessity for temporal logic annotations and highlight the difficulty models face when reasoning across video frames. In parallel, spatial datasets such as SpatialVLM (Chen et al., 2024a), WhatsUp (Kamath et al., 2023), DC2 (Wang et al., 2024e), and Grounded (Chen et al., 2024f) showcase improvements through extensive spatial reasoning QA pairs and challenging spatial configurations.
Iterative and Reflective Reasoning Iterative reasoning capabilities have markedly improved through datasets like VISCO (Wu et al., 2024b), Mulberry-260k (Yao et al., 2024), FIRE (Li et al., 2024d), VLFeedback (Li et al., 2024c), Silkie (Li et al., 2023b), TIIL (Huang et al., 2024c), ConMe (Huang et al., 2024a), and Reflective (Zhang et al., 2024f), enabling models to learn from mistakes by both incorporating reflective and iterative feedback loops. For instance, VISCO demonstrated that human critiques significantly outperform model critiques (76% error correction), highlighting a persistent gap in models’ ability to independently self-correct.
Complex Multimodal Evaluation Complex multimodal datasets like EMMA (Hao et al., 2025), CoMT (Cheng et al., 2024c), SMIR (Li et al., 2025a), ProVision (Zhang et al., 2024e), MAGEBench (Zhang et al., 2024i), MM-Vet v2 (Yu et al., 2024), JourneyBench (Wang et al., 2025), VERIFY (Bi et al., 2025) and CompCap (Chen et al., 2024g) exposed limitations in current multimodal models’ integration capabilities. EMMA specifically shows that models struggle significantly with iterative multimodal interactions, particularly when tasks demand deep integration across domain-specific problems. Benchmarks like FineCops-Ref (Liu et al., 2025b), VL-RewardBench (Li et al., 2024b), MM-SAP (Wang et al., 2024h), PCA-Bench (Chen et al., 2024d), AttCoSeg (Pramanick et al., 2024), and M3CoT (Chen et al., 2024e) emphasize detailed evaluation metrics (Recall@k, AUROC). These benchmarks provide insights into models’ actual reasoning capabilities beyond traditional accuracy metrics.
Counterfactual and Logical Reasoning Datasets like CounterCurate (Zhang et al., 2024c), C-VQA (Zhang et al., 2024g), CausalChaos! (Parmar et al., 2024), LogicAI (Xiao et al., 2024), and MCDGRAPH (Zhu et al., 2024) explicitly test model reasoning robustness, showing improvements through challenging counterfactual examples and structured logical questions.
Active Perception and Progressive Reasoning Active perception datasets such as ActiView (Wang et al., 2024j) and progressive reasoning benchmarks such as Blink (Fu et al., 2024b), ADL-X (Reilly et al., 2024), and GUIDE (Liang et al., 2024a) challenge models’ dynamic reasoning capabilities through active view adjustments and frame-level progression.
6 Conclusion
In this paper, we presented a comprehensive survey of existing work on reasoning in Multimodal Large Language Models (MLLMs) and Large Language Models (LLMs), with a focus on optimizing model performance and identifying the most effective reasoning trajectories. We introduced a variety of post-training and test-time computation methods, discussing their potential to enhance model capabilities on complex, real-world tasks. Our analysis highlights the critical role of reasoning in improving the visual understanding abilities of MLLMs. Furthermore, we have provided insights and outlined potential directions for future research to guide the development of more robust and efficient reasoning frameworks.
Appendix A Implications from the Learning Science Perspective
The design of CoT reasoning for LLMs aligns with several key cognitive science and learning science theories that enhance structured reasoning, problem-solving, and knowledge construction.
One key theory is Cognitive Load Theory (Sweller, 2011; Paas et al., 2004), which reduces a human’s working memory workload by breaking complex problems into manageable steps. CoT reasoning follows this approach by prompting models to generate step-by-step solutions rather than arriving at an answer in one leap. This structured decomposition mirrors how human learners handle intricate problems by offloading cognitive effort across multiple processing stages. Similarly, in the context of learning, scaffolding refers to offering temporary assistance to learners, enabling them to accomplish tasks that would be difficult or impossible for them to complete independently. This support is gradually removed as learners develop the skills and confidence required to perform the task on their own (Van de Pol et al., 2010). This supports CoT’s design by guiding LLMs through intermediate steps, much like how educators provide structured support to help students grasp new concepts before transitioning to independent problem-solving. These theories are both rooted in the Zone of Proximal Development (ZPD), a key construct in Lev Vygotsky’s theory of learning and development (Vygotsky, 1978).
CoT reasoning also aligns with metacognition (Lai, 2011) and self-explanation (Bisra et al., 2018; Chi et al., 1994) by encouraging models to “think about their thinking”. Just as self-explanation improves human learning by prompting individuals to articulate their reasoning (Chi et al., 1994; VanLehn et al., 1992), CoT forces LLMs to justify their steps, reducing reliance on shallow heuristics. This ties into Dual-Process Theory (Frankish, 2010), where CoT shifts LLMs from fast, intuitive decision-making (System 1) to deliberate, analytical reasoning (System 2), leading to more logical and consistent outputs.
Moreover, CoT fosters constructivist learning (Waite-Stupiansky, 2022; Bruner, 2009) by enabling LLMs to incrementally build knowledge structures through active reasoning. Instead of passively retrieving answers from training data, the model synthesizes prior knowledge with new information, improving adaptability. Additionally, analogical reasoning (Gentner & Maravilla, 2017) plays a role in CoT by helping LLMs map relationships between concepts, allowing them to generalize problem-solving strategies across different contexts.
By integrating these cognitive science and learning science principles, CoT reasoning enhances the interpretability, reliability, and generalization of LLM outputs, making them more aligned with human cognitive processes and educational best practices.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=10em,font=,, where level=3text width=8.2em,font=,, [ Inference [ Adaptive
Inference [  Guo et al. (2025),  Hu et al. (2025),  Yang et al. (2024b),
 Xue et al. (2024),  Wu et al. (2024b),  Wang et al. (2024f),
 Zou et al. (2024a),  Qu et al. (2024),  Zhang et al. (2024k),
 Huo et al. (2024),  Liao et al. (2024b),  Su et al. (2024),
 Li et al. (2024e),  Wang et al. (2024e),  Xu et al. (2025),
 Zhao et al. (2024),  Liao et al. (2024a),  Mahmood et al. (2024),
 Cao et al. (2024b),  Wu et al. (2024a),  Yang et al. (2024a),
 Li et al. (2023a),  Pi et al. (2023) , leaf, text width=30em ] ] [ Reward
Models [  Guo et al. (2025),  Hao et al. (2025),  Li et al. (2024b),
 Fu et al. (2024c),  Ke et al. (2024),  Xiong et al. (2024),
 Cascante-Bonilla et al. (2024),  Cheng et al. (2024b),
 Zhang et al. (2024k), , leaf, text width=30em ] ] [ Feedback [  Guo et al. (2025),  Wu et al. (2024b),  Hu et al. (2025),
 Meng et al. (2025),  Li et al. (2024d),  Xiong et al. (2024),
 Cheng et al. (2024b),  Zhang et al. (2024k),  Zheng et al. (2024),
 Liao et al. (2024a),  Han et al. (2024),  Wang et al. (2024i),
 Shao et al. (2024),  Liang et al. (2024b),
 Lee et al. (2024),  Hu et al. (2024), , leaf, text width=30em ] ] [ Multimodal
Decomposition [  Xiang et al. (2025),  Liang et al. (2025),  Liu et al. (2025a),
 Shlapentokh-Rothman et al. (2024),  Zhang et al. (2024b),
 Chen et al. (2024f),  Hsu et al. (2024),  Su et al. (2024),
 Reilly et al. (2024),  Wang et al. (2024a),  Zhang et al. (2024h),
 Min et al. (2024),  Parmar et al. (2024),
 Wang et al. (2024b),  Dzabraev et al. (2024),  Wu et al. (2024a),
 Wu & Xie (2023),  Zhang et al. (2023), , leaf, text width=30em ] ] [ Efficiency and
Scalability [  Goyal et al. (2025),  Li et al. (2025b),  Sun et al. (2024),
 Fan et al. (2024),  Park et al. (2025),  Wang et al. (2024e),
 Wang et al. (2025),  Huo et al. (2024),
 Lin et al. (2024),  Ni et al. (2024),  Wang et al. (2024i),
 Huang et al. (2024a),  Zhang et al. (2024m),  Zhan et al. (2024),
 Chen et al. (2024b),  Wang et al. (2024b),  Zhao et al. (2023),
 Pi et al. (2023),  Chen et al. (2024a) , leaf, text width=30em ] ] [ Search
Strategies [  Yao et al. (2024),  Yang et al. (2024b),  Wang et al. (2024f),
 Imam et al. (2025),  Xu et al. (2025),  Qu et al. (2024),
 Wang et al. (2024e),  Li et al. (2024e),  Kazemi et al. (2024),
 Shin et al. (2024),  Zhang et al. (2024l),
 Zhao et al. (2024),  Fu et al. (2024c),  Mahmood et al. (2024),
 Wu & Xie (2023),  Hu et al. (2023), , leaf, text width=30em ] ] ]
Figure 4: Comprehensive Overview of Methods and Frameworks focus on test-time compute
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=10em,font=,, where level=3text width=8.2em,font=,, [ Post
Training [ Policy
Optimization [  Fu et al. (2024c),  Zhang et al. (2024d),  Hu et al. (2024),
 Tian et al. (2024),  Kim et al. (2023),  Yan et al. (2024),
 Ouali et al. (2024),  Xiong et al. (2024),  Wang et al. (2024j),
 Waite et al. (2025),  Wei et al. (2025),  Zhao et al. (2025),
 Xu et al. (2025),  Deng et al. (2024), , leaf, text width=30em ] ] [ Model
Architecture [  Reilly et al. (2024),  Ma et al. (2024a),  Zhang et al. (2024j),
 Zhan et al. (2024),  Jiao et al. (2024),
 He et al. (2024),  Ma et al. (2023),  Ben-Kish et al. (2024),
 Hu et al. (2024),  Wang et al. (2024d),  Wang et al. (2024c),
 Cao et al. (2024a),  Chen et al. (2024f),  Huang et al. (2024d), , leaf, text width=30em ] ] [ Reward
Alignment [  Hao et al. (2025),  Li et al. (2024b),  Xiong et al. (2024),
 Ke et al. (2024),  Fu et al. (2024c),  Dzabraev et al. (2024),
 Hu et al. (2024), , leaf, text width=30em ] ] [ Iterative
Refinement [  Meng et al. (2025),  Zheng et al. (2024),  Liao et al. (2024a),
 Wu et al. (2024b),  Shao et al. (2024),
 Han et al. (2024),  Hu et al. (2024),  Lee et al. (2024),
, leaf, text width=30em ] ] [ Spatial
Temporal
Modeling [  Wang et al. (2024i),  Zhang et al. (2024m),  Chen et al. (2024a),
 Wu & Xie (2023), , leaf, text width=30em ] ] [ Dataset
Curation [  Fan et al. (2024),  Wang et al. (2025),
 Nagarajan & Torresani (2024),  Zhang et al. (2024l),
 Lin et al. (2024),  Zhao et al. (2023),  Hu et al. (2023), , leaf, text width=30em ] ] ]
Figure 5: Comprehensive Overview of Methods and Frameworks focus on post-training improvement