### v2

# Why Reasoning Matters? A Survey of Advancements in Multimodal Reasoning (v1)

Abstract
Reasoning is central to human intelligence, enabling structured problem-solving across diverse tasks. Recent advances in large language models (LLMs) have greatly enhanced their reasoning abilities in arithmetic, commonsense, and symbolic domains. However, effectively extending these capabilities into multimodal contexts—where models must integrate both visual and textual inputs—continues to be a significant challenge. Multimodal reasoning introduces complexities, such as handling conflicting information across modalities, which require models to adopt advanced interpretative strategies. Addressing these challenges involves not only sophisticated algorithms but also robust methodologies for evaluating reasoning accuracy and coherence. This paper offers a concise yet insightful overview of reasoning techniques in both textual and multimodal LLMs. Through a thorough and up-to-date comparison, we clearly formulate core reasoning challenges and opportunities, highlighting practical methods for post-training optimization and test-time inference. Our work provides valuable insights and guidance, bridging theoretical frameworks and practical implementations, and sets clear directions for future research.
1 Introduction
Reasoning is a fundamental aspect of human intelligence, enabling us to solve complex problems effectively. In LLMs, reasoning has shown promising capabilities, leading to significant advancements in various domains (Bi_2024), including arithmetic, commonsense, and symbolic reasoning. For instance, Chain-of-Thought (CoT) prompting mimics human stepwise problem-solving to boost LLM performance, especially in very large models.
Beyond CoT, other prompting strategies, such as generated knowledge prompting and tree search algorithms (e.g., STAR search (2203.14465-zelikman_star:_2022)), have been incorporated into LLM training to further enhance reasoning capabilities. These methods encourage the model to articulate its internal reasoning in natural language, which reinforces its understanding of the task. By laying out intermediate steps, the model can verify and adjust its logic before arriving at the final answer.
Figure 1: Papers on visual reasoning per quarter over the last three years, with state computed using referenced papers. (Data current to mid-March 2025.)
While reasoning in LLMs has progressed significantly, extending these capabilities to multimodal tasks remains an emerging frontier, as multimodal large language models (MLLMs) must navigate the added complexity of interpreting and integrating information from both visual and textual modalities, often resolving ambiguities, inconsistencies, or gaps that arise when the two sources conflict or diverge.
Effective reasoning is central to enabling MLLMs to achieve a compositional understanding—allowing them to deconstruct complex tasks into interpretable, modality-spanning steps. It also supports iterative processes such as error correction and self-refinement, where models can revise their outputs by reevaluating both visual and textual cues.
Moreover, strong reasoning capabilities help clear up confusion by inferring spatial relationships, handling counterfactuals, and choosing appropriate tools or actions. These mechanisms also play a crucial role in mitigating hallucinations by grounding model outputs in cross-modal evidence, ultimately improving both accuracy and trustworthiness. Finally, reasoning extends the potential of MLLMs to handle hypothetical scenarios, helping them look ahead and explore “what-if” situations that span both language and vision.
Initial research has demonstrated that integrating Chain-of-Thought (CoT) reasoning across different modalities can significantly enhance the performance of MLLMs. For example, 2412.18319-yao_mulberry:_2024; 2410.02884-zhang_llama-berry:_2024; bi2024unveilingvisualperceptionlanguage reports that the LLama series models’ reasoning abilities are greatly improved by solely utilizing filtered self-generated reasoning paths.
Building on this momentum, our work provides a comprehensive overview of reasoning in both LLMs and MLLMs, with a focus on its techniques, applications, and future directions. We aim to bridge the gap between text-based and multimodal reasoning, offering insights into how reasoning can further enhance the capabilities of LLMs in multimodal contexts. In the following sections, we begin with a clear problem formulation and definition of reasoning, followed by an in-depth analysis of reasoning techniques—from post-training strategies to test-time computation. We then examine recent trends in datasets and benchmarking. Our goal is to offer a well-organized and accessible survey that supports both theoretical understanding and practical implementation of reasoning in LLMs and MLLMs.
2 Background
In complex question-answering tasks, directly predicting an answer can be highly uncertain due to the vast range of possible responses. A more effective approach involves breaking down the reasoning process into a sequence of intermediate steps. This structured method not only improves interpretability but also helps reduce uncertainty at each inference step.
Formally, let $Q$ denote a given question. Conventional language models typically aim to model the conditional probability $P(A\mid Q)$ of the answer $A$ given the question. However, as the complexity of $Q$ increases, the prediction becomes more uncertain—reflected by a rise in the conditional entropy $H(A\mid Q)$. A natural approach to mitigate this uncertainty is to decompose the reasoning process into a sequence of intermediate steps. This leads to modeling the answer generation as a structured chain of conditional probabilities:
$P(\text{Step}_{1}\mid Q)\cdot P(\text{Step}_{2}\mid\text{Step}_{1},Q)\cdot\dots\cdot P(\text{Step}_{t}\mid\text{Step}_{1:t-1},Q)\cdot P(A\mid\text{Step}_{1:t},Q)$ (1)
Such a decomposition encourages the model to reduce uncertainty at each stage and improves interpretability and robustness in complex question answering. This step decomposition is useful because conditioning on prior steps reduces uncertainty. Specifically, the conditional entropy satisfies: $H(\text{Step}_{t}\mid\text{Step}_{<t},Q)\leq H(\text{Step}_{t}\mid Q)$, meaning each step becomes easier to predict as more context accumulates. From a probabilistic standpoint, the entire reasoning process can be viewed as a trajectory, denoted as $\tau=(\text{Step}_{1},\dots,\text{Step}_{t})$, which represents a complete reasoning path. Using this, the overall probability of an answer given a question can be written as a marginal over all possible trajectories:
$P(A\mid Q)=\sum_{\tau}P(A\mid\tau,Q)\cdot P(\tau\mid Q).$ (2)
However, this formulation presents a challenge: enumerating all possible reasoning paths $\tau$ is intractable. In practice, CoT reasoning offers a practical workaround by sampling a single trajectory $\hat{\tau}$ from the distribution $P(\tau\mid Q)$, and using it to approximate the answer:
$P(A\mid Q)\approx P(A\mid\hat{\tau},Q),\quad\text{where}\quad\hat{\tau}\sim P(\tau\mid Q).$ (3)
This is beneficial for many tasks if the model is capable of generating one coherent reasoning path that captures the correct structure of the solution, thereby reducing uncertainty and improving answer accuracy. More advanced methods such as Tree-of-Thought (ToT) expand on CoT by generating and evaluating multiple reasoning paths, and can be seen as a multiple point estimation approach. In this case, the model generates a set of reasoning paths $\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ and evaluates them to select the most promising one.
Given the challenge of enumerating all possible reasoning paths, our objective becomes clear: to generate and identify only the most promising trajectories. Achieving this effectively requires two key components: (i) Generation – a strong model capable of producing high-quality reasoning paths $\tau$; (ii) Selection – a reliable mechanism for evaluating and choosing the best candidate path.
These components are deeply interconnected: improved generation policies yield better trajectories, which in turn provide stronger supervision signals to further refine the model. This creates a virtuous cycle of mutual enhancement, as illustrated in Figure 2.
Figure 2: Framework illustrating training and inference for reasoning optimization. A virtuous cycle emerges as better policies generate improved trajectories, which in turn enhance the model through stronger supervision.
At inference time, we want to search for the best $\tau$ under the current model. During training, we aim to improve the model so that it generates reasoning paths that are more likely to lead to correct answers.
Broadly speaking, recent approaches fall into two main categories:
Post-training improvement involves optimizing the model policy $\pi$ to better align outputs with desired utility, typically through techniques like fine-tuning or reinforcement learning. Formally, this corresponds to solving Equation (4), where the goal is to learn a policy that maximizes expected utility over generated outputs $\tau$.
Test-time compute focuses on improving LLM performance during inference without modifying the model’s core parameters. This aligns with the objective in Equation (4), where the goal is to select the most useful output trajectory $\tau$ under a fixed policy $\pi^{*}$, optimizing accuracy, coherence, and efficiency through dynamic strategies like CoT, ToT, and search-based reasoning.
$\begin{cases}\pi^{*}=\displaystyle\arg\max_{\pi}\,\mathbb{E}_{\tau\sim\pi}[\text{Utility}(\tau)]&\text{(Training: Learn a better model)}\\ \tau^{*}=\displaystyle\arg\max_{\tau\sim\pi^{*}}\,\text{Utility}(\tau)&\text{(Inference: Select best trajectory)}\end{cases}$ (4)
In the following sections, we will discuss these two components in detail.
3 Post-training improvements
To enhance the quality of reasoning, we aim to learn a model that generates high-utility reasoning trajectories. One effective approach frames this as a Markov Decision Process (MDP), where the model is trained to maximize the expected return of a reasoning path $\tau$:
$\max_{\theta}\ \mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t}\gamma^{t}R(s_{t},a_{t})\right]\quad\text{with}\quad\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})A(s_{t},a_{t})\right]$ (5)
$s_{t}$ is the sequence of tokens up to time $t$, and $a_{t}$ is the next token. The advantage function $A(s_{t},a_{t})$ estimates the benefit of choosing $a_{t}$ in state $s_{t}$, guiding the model toward higher-utility token sequences. This objective can be directly optimized with respect to the model parameters using policy gradient methods (williams1992simple). Notably, this formulation enables the training of LLMs not merely to predict the next token, but to generate complete reasoning trajectories that are optimized for long-term rewards.
ȷ Policy optimization is still in its early stages, often relying on existing methods with a primary focus on text token optimization. However, approaches from visual reinforcement learning are worth exploring to better align policies with perceptual and spatial dynamics. Reward alignment remains heavily reliant on textual signals, highlighting an opportunity to shift focus toward visual-based rewards. Developing methods that can interpret and adapt to visual feedback—such as spatial cues, object dynamics, and scene changes—will be key to achieving more grounded and effective visual decision-making. In terms of model architecture, improvements often stem from better attention to visual details that align with long-horizon reasoning steps. Spatial-temporal modeling offers unique opportunities for more creative action definition to manipulate the visual information for better reasoning. Similarly, during data curation, verifiers commonly depend on grounding models or language cues—indicating potential for more creative, vision-focused verifier designs.
3.1 Policy Optimization
To solve the above reward maximization problem, recent advancements in visual reasoning have prominently focused on integrating reinforcement learning (RL) and imitation learning (IL) to align MLLMs more closely with human reasoning. IL approaches, such as Thought Cloning (2503.08525-wei_gtr:_2025), surpass traditional cloning by aligning intermediate reasoning steps instead of final outputs alone. Coupled with iterative methods like DAgger, this reduces dataset shift and hallucination. LLaVA-Critic (2410.02712-xiong_llava-critic:_2024) further generalizes reward signals via preference-based alignment, boosting multimodal reasoning effectiveness with minimal modality-specific fine-tuning. RL from Simulations (RLS3) (2501.18880-waite_rls3:_2025), employing Soft Actor-Critic (SAC) agents, significantly enhances spatial reasoning accuracy and efficiency in models like PaliGemma, outperforming random exploration baselines (2501.18880-waite_rls3:_2025). FuRL (2406.00645-fu_furl:_2024) addresses RL reward misalignment through dual-stage alignment, effectively mitigating sparse reward challenges. Adaptive Reward Programming (ARP-DT+) (2309.10790-kim_guide_2023), integrating CLIP representations with Decision Transformers, enhances reasoning robustness across domains by employing active inference to counter distractions and domain shifts.
3.2 Reward Alignment
Recent approaches have explored iterative refinement and reflection mechanisms, token-level rewards, and automated benchmarking to enhance reasoning quality and mitigate hallucinations. PARM++ (2501.13926-guo_can_2025) introduces a reflection-based mechanism that iteratively refines outputs by identifying misalignments and actively correcting them through self-assessment loops, significantly improving test-time GenEval scores up to three reflection cycles (2501.13926-guo_can_2025). Similarly, REVERIE employs rationale-based training, explicitly supervising reasoning steps, thus improving coherence and reducing hallucinations compared to models without rationales (2407.11422-zhang_reflective_2024). Fine-grained reward models, such as MMViG (2402.06118-yan_vigor:_2024) and Reward Alignment via Preference Learning (RAPL) (2310.07932-tian_what_2024), emphasize granular feedback at individual reasoning steps, achieving precise improvements without manual reward tuning. Similarly, CLIP-DPO (2408.10433-ouali_clip-dpo:_2024) leverages pre-trained CLIP directly as a reward function, streamlining reward modeling. Challenges such as constrained perceptual fields are addressed through dynamic evaluation strategies like zooming and shifting, emphasizing the importance of sequential decision-making in visual reasoning (2410.04659-wang_actiview:_2024). Structural issues like positional bias and patch-boundary effects further underscore the need for targeted architectural optimizations (2402.07384-zhang_exploring_2024). Recent approaches integrate reward signals directly into generation. TLDR (2410.04734-fu_tldr:_2024) applies token-level binary rewards, enabling real-time feedback to mitigate hallucinations and boost both annotation efficiency and backbone model performance (2410.04734-fu_tldr:_2024). Complementarily, EACO enhances intermediate object recognition, significantly reducing hallucinations compared to earlier models like LLaVA-v1.6-7B (2412.04903-wang_eaco:_2024). Distillation frameworks also contribute: FIRE leverages iterative student–teacher feedback and structured evaluation to refine visual reasoning (2407.11522-li_fire:_2024), while SILKIE demonstrates the scalability of GPT-4V-generated feedback in improving perception and cognition without human intervention (2312.10665-li_silkie:_2023).
3.3 Model Architecture
Nested architectures, such as MaGNeTS, have emerged as effective means of balancing computational efficiency with model accuracy by employing parameter-sharing and caching mechanisms during inference (2502.00382-goyal_masked_2025). Intrinsic activation approaches, notably ROSS, bypass external modules by integrating multimodal understanding directly into the model’s core, yielding improved adaptability to diverse visual inputs, including depth maps (2410.09575-wang_reconstructive_2024). Additionally, DIFFLMM incorporates diffusion models and attention-based segmentation, enhancing visual grounding precision (2410.08209-cao_emerging_2024), while Mini-Monkey targets the ”sawtooth effect” in lightweight models through adaptive cropping and scale compression, achieving superior visual-text alignment without extensive computational resources (2408.02034-huang_mini-monkey:_2024). SEA dynamically adjusts embedding alignment strategies to maintain performance across varying resolutions and model sizes (2408.11813-yin_sea:_2024). Concurrently, PAE-LaMM demonstrates the combined effectiveness of vision encoder adaptation and pixel reconstruction tasks, systematically enhancing visual detail awareness and question-answering performance (2408.03940-gou_how_2024). Conversely, LLAVIDAL and EventLens uniquely integrate domain-specific visual elements, such as object-centric embeddings and event-aware tokens, respectively, directly within LLM architectures to effectively reason about activities and temporal contexts (2406.09390-reilly_llavidal:_2024; 2404.13847-ma_eventlens:_2024).
3.4 Spatial Temporal Modeling
Unlike pure language reasoning, visual reasoning involves a richer action space, allowing operations such as zooming, cropping, and frame selection, which enables the model to interact with and manipulate visual inputs. The GeoGLIP pipeline utilizes geometric pre-training with a dynamic feature router to advance fine-grained visual mathematical understanding by balancing symbolic and visual modalities based on task demands (2501.06430-zhang_open_2025). Concurrently, hallucination issues are mitigated via targeted fine-tuning and prompting strategies grounded in recent vision-language work like CapQA (2501.02964-hu_socratic_2025). For spatial reasoning, simpler graph topologies have been shown to enhance performance, while structural complexity reduces accuracy in edge-centric tasks (2412.13540-zhu_benchmarking_2024). Augmentation via uncertainty-aware active inference further boosts attribute detection precision (2412.07012-zhang_provision:_2024). Intrinsic spatial-temporal modeling is advanced through architectures embedding multi-modal understanding without reliance on external depth tools (2410.09575-wang_reconstructive_2024), and dual-branch structures improve video temporal grounding in multi-hop reasoning (2408.14469-chen_grounded_2024). The shift toward multimodal benchmarks better reflects real-world spatial-temporal reasoning (2408.08632-li_survey_2024). Integrating rationalization with answer generation enhances temporal reasoning (2407.06189-zohar_video-star:_2024), while automated annotation improves data scalability and bridges gaps between commercial and open-source Video-LLMs (2406.11303-li_videovista:_2024). Multi-modal integration and active inference optimize modality selection for multi-step spatial-temporal reasoning (2405.03272-zhang_worldqa:_2024). Adjusting learning objectives to minimize hallucinations improves precision and EOS decision-making (2402.14545-yue_less_2024).
Figure 3: Search framework where language models explore and refine reasoning paths. Trajectories are scored using reward models, based on expected utility or final output quality, and guided by feedback, world models, and evaluators to select the most promising steps.
4 Test-Time compute
At test time, the goal is to find the reasoning trajectory $\tau$ that maximizes a utility function as shown as red and green path in Figure 3
$\tau^{*}=\arg\max_{\tau}\,\mathcal{U}(\tau),\quad\text{where }\mathcal{U}(\tau)=\begin{cases}\sum_{t}r(s_{t},a_{t})&\text{(MDP-style reward)}\\ \mathbb{P}(\text{goal}\mid\tau)&\text{(Goal likelihood)}\\ \mathbb{P}(\tau\succ\tau^{\prime})&\text{(Preference-based)}\\ f(\text{rank}(\tau))&\text{(Ranking-based)}\\ \vdots&\text{(Others: risk-sensitive, etc.)}\end{cases}$ (6)
The utility $\mathcal{U}(\tau)$ can be defined in various ways, including cumulative rewards (MDP-style), goal likelihood, preference comparisons, and ranking-based scores. The choice of utility depends on the supervision available and the nature of the task. The core challenge remains sampling high-utility trajectories under the chosen formulation.
Since directly modeling $\mathcal{U}(\tau)$ is often difficult, we use reward models to approximate it. These models can include (1) absolute rewards (as in standard RL), (2) pairwise preference models (e.g., DPO), and (3) ranking-based models.
ȷ Search Strategies Current search algorithms primarily operate on distinct textual tokens. However, innovative approaches are emerging for raw visual tokens, termed Creative Visual Search, expanding search capabilities directly within visual spaces. Moreover, Multi-Granularity Spatial Search is gaining attention as an effective extension, enhancing exploration across diverse spatial granularities. Reward and Feedback Reward systems can significantly benefit from deeper visual insights. Leveraging visual prompts encourages models to engage in self-reflection grounded in visual information, substantially improving semantic alignment and decision-making effectiveness. Iterative Refinement Determining the alignment between textual tokens and visual information remains challenging. Consequently, model self-reflection can produce hallucinated interpretations. Nevertheless, these reflections offer valuable insights, highlighting new pathways to mitigate unnecessary or hallucinated reflections and enhance model reliability.
4.1 Search Strategies
Monte Carlo Tree Search (MCTS) has emerged as a powerful tool for managing uncertainty and sequential decision-making. 2501.10674-imam_can_2025 applied MCTS to enhance temporal reasoning, reducing hallucinations in time-dependent data. Building on this, CoMCTS(2412.18319-yao_mulberry:_2024) unified multiple MLLMs within a collaborative tree search, addressing model bias and improving accuracy and efficiency. Vision-specific strategies advanced with VisVM (2412.03704-wang_scaling_2024), which estimates long-term candidate value to reduce hallucinations, outperforming immediate-alignment methods like CLIP-PRM. Video reasoning similarly benefited from reward modeling and adaptive exploration to prioritize key frames (2412.10471-yang_vca:_2024). Structured search methods also gained traction. LLaVA-o1 (2411.10440-xu_llava-cot:_2025) uses stage-level beam search for complex reasoning, while MVP (2408.17150-qu_look_2024) aggregates certainty across multi-perspective captions to resist adversarial inputs. DC2 (2408.15556-wang_divide_2024) applies MCTS-based cropping to focus on salient image regions for high-resolution reasoning. Multimodal and temporal search frameworks like VideoVista (2406.11303-li_videovista:_2024), WorldRetriver (2405.03272-zhang_worldqa:_2024), and DynRefer (2405.16071-zhao_dynrefer:_2024) surpass static baselines using adaptive sampling, fusion, and stochastic inference. Step-by-step comparative reasoning with LMs also enhances video QA (2404.16222-nagarajan_step_2024). Innovations in context refinement (VURF (2403.14743-mahmood_vurf:_2024)), image decomposition (V*(2312.14135-wu_v*:_2023)), and dynamic tool use (AVIS(2306.08129-hu_avis:_2023)) highlight the shift toward adaptive visual reasoning. FuRL (2406.00645-fu_furl:_2024) aligns reward modeling with iterative fine-tuning for better performance. In embodied AI, combining chain-of-thought, self-verification, and MCTS-driven planning enables scalable, robust decision-making in dynamic environments (2404.15190-shin_socratic_2024).
4.2 Adaptive Inference
Adaptive inference is reshaping vision-language reasoning by enabling dynamic, context-sensitive processing that improves both accuracy and efficiency. Central to this is the iterative evaluation and refinement of outputs, often through internal feedback and external verification. LOOKBACK enhances correction accuracy through iterative visual re-examination within each reasoning step (2412.02172-wu_visco:_2024). IXC-2.5 balances performance and response length via Best-of-N sampling guided by reward models (2501.12368-zang_internlm-xcomposer2.5-reward:_2025), while PARM++ uses reflection-based refinement to align generated images with prompts (2501.13926-guo_can_2025). Further innovations include ProgressCaptioner’s sliding-window approach for tracking action progress over time (2412.02071-xue_progress-aware_2024). To combat multi-modal hallucinations, MEMVR triggers visual retracing based on uncertainty, optimizing cost-accuracy tradeoffs (2410.03577-zou_look_2024), while MVP aggregates certainty across diverse views (2408.17150-qu_look_2024). SID applies token-level contrastive decoding to filter irrelevant content (2408.02032-huo_self-introspective_2024). Models like DualFocus integrate macro and micro reasoning for fine-grained attention control (2402.14767-cao_dualfocus:_2024), and LISA++ enables test-time compute scaling without retraining (2312.17240-yang_lisa++:_2024). PerceptionGPT accelerates inference via adaptive token embeddings that encode spatial cues and dynamically weigh layers (2311.06612-pi_perceptiongpt:_2023).
4.3 Reward
2501.13926-guo_can_2025 introduce ORM and PRM, showing that ORM’s final-step evaluation significantly enhances image generation through ranking data. PARM++ further refines this method by enabling iterative reflection for self-correction, thus improving prompt fidelity. Complementarily, 2501.05444-hao_can_2025 highlight persistent errors in VLM spatial reasoning and explanation, proposing reward re-ranking as an interim solution. To assess VL-GenRMs more comprehensively, 2411.17451-li_vlrewardbench:_2024 propose VL-RewardBench, pinpointing perception and reasoning challenges while advocating critic training and co-evolutionary reward learning. Self-training methods also gain prominence: R3V (2411.00855-cheng_vision-language_2024) leverages synthesized rationales to enhance noise-robust Chain-of-Thought (CoT) refinement, while CECE (2410.22315-cascante-bonilla_natural_2024) enriches evaluation using LLM-generated entailment/contradiction captions. Enhanced supervision methods also emerge. 2410.16198-zhang_improve_2024 enhance CoT reasoning through ShareGPT-4O distillation, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). LLaVA-Critic (2410.02712-xiong_llava-critic:_2024) builds upon this using critic-guided DPO without external human feedback. FuRL (2406.00645-fu_furl:_2024) introduces a two-stage tuning approach addressing misalignments and temporal errors, while HYDRA (2403.12884-ke_hydra:_2024) employs dynamic loops for adaptive instruction ranking. Additionally, m&m’s (2403.11085-ma_m&ms:_2024) provides a structured framework for evaluating multi-step planning via tool and argument accuracy.
4.4 Feedback
VOLCANO employs a three-stage iterative framework—initial response generation, visual self-assessment, and revision—to reduce hallucinations by emphasizing accurate image details (2311.07362-lee_volcano:_2024). Similarly, LOOKBACK mandates explicit atomic verification against images to enhance both critique and correction processes for LVLMs (2412.02172-wu_visco:_2024). LLaVA-Critic advances feedback by incorporating diverse critic instructions and iterative DPO training, leveraging internal data (2410.02712-xiong_llava-critic:_2024). Additionally, tailored visual prompting with iterative binary verification effectively enhances semantic grounding in models like LLaVA-1.5, ViP-LLaVA, and CogVLM (2404.06510-liao_can_2024). Minimalist reinforcement learning setups have shown effectiveness, particularly for multimodal mathematical reasoning, with larger models like InternVL2.5-38B benefiting from difficulty-based data filtering to stabilize training (2503.07365-meng_mm-eureka:_2025). Explicit visual CoT mechanisms, exemplified by VisCoT, integrate bounding box predictions and iterative cropping to enhance visual question-answering performance (2403.16999-shao_visual_2024). In video understanding, VIDEOTREE utilizes adaptive hierarchical clustering for efficient navigation through structured tree-based representations (2405.19209-wang_videotree:_2024). Lastly, prompt engineering strategies like MiHO and MiHI significantly reduce hallucinations without model retraining, post-processing evaluations with GPT-4 further enhancing reliability without architectural modifications (2402.01345-han_skip_2024; 2501.02964-hu_socratic_2025).
4.5 Iterative Refinement
Adaptive inference transforms vision-language reasoning by facilitating dynamic, context-sensitive processing to enhance accuracy and efficiency. Central to this approach is iterative output refinement through internal feedback and external verification. For instance, IXC-2.5 employs Best-of-N sampling guided by reward models to optimize performance and response length (2501.12368-zang_internlm-xcomposer2.5-reward:_2025), while PARM++ uses reflection-based refinement for better alignment between images and prompts (2501.13926-guo_can_2025). Further advancements include LLaVA-o1’s structured perception and stage-level beam search for decomposing complex reasoning tasks (2411.10440-xu_llava-cot:_2025). To mitigate multi-modal hallucinations, MEMVR implements visual retracing based on uncertainty to optimize the cost-accuracy tradeoff (2410.03577-zou_look_2024), MVP aggregates certainty across diverse views (2408.17150-qu_look_2024), and SID employs token-level contrastive decoding to eliminate irrelevant content (2408.02032-huo_self-introspective_2024). Moreover, DualFocus integrates macro- and micro-level reasoning for precise attention control (2402.14767-cao_dualfocus:_2024), LISA++ enables test-time compute scaling without retraining (2312.17240-yang_lisa++:_2024), and PerceptionGPT accelerates inference through adaptive token embeddings encoding spatial cues and dynamically weighted layers (2311.06612-pi_perceptiongpt:_2023).
4.6 Dataset Curation
Recent work shows that high-quality, strategically curated datasets outperform large but noisy alternatives in guiding reasoning paths. R-CoT (2410.17885-deng_r-cot:_2024) enhances geometric reasoning via reverse generation and stepwise synthesis. Task-specific datasets such as SMIR (2501.03675-li_smir:_2025) show up to 8% performance gains over generic datasets, demonstrating the value of targeted curation. Gradual complexity in dataset construction, as shown in VIREO (2406.19934-cheng_least_2024), ensures foundational reasoning skills before introducing advanced tasks. Scalable dataset creation through automation also proves effective. VideoVista (2406.11303-li_videovista:_2024) uses auto-annotations for video reasoning, outperforming manually curated sets. DecoVQA+ (2409.19339-zhang_visual_2024) explicitly teaches when to apply question decomposition, aided by SelectiveVQD loss for better strategy selection. LocVLM (2404.07449-ranasinghe_learning_2024) scales with pseudo-data and implicit feedback signals like location and relevance prediction. CogCoM (2402.04236-qi_cogcom:_2024) incorporates tasks like grounding and manipulation to reinforce step-by-step reasoning.
5 Datasets and Benchmarks
Structured and Task-Specific Reasoning Datasets such as Visual-RFT(2503.01785-liu_visual-rft:_2025), CapQA(2501.02964-hu_socratic_2025), GUIDE(2406.18227-liang_guide:_2024), STAR(2405.09711-wu_star:_2024), Visual Genome(2407.11422-zhang_reflective_2024), VL-GPT(2312.09251-zhu_vl-gpt:_2023), and Interfacing (2312.07532-zou_interfacing_2024) have introduced structured prompting that significantly enhances models’ reasoning accuracy by explicitly guiding reasoning processes. In contrast to earlier benchmarks like STAR, these datasets place a stronger emphasis on task-specific reward functions and structured inference, aiming to improve generalization across diverse visual and linguistic contexts.
Temporal and Spatial Reasoning Temporal benchmarks like VisualQA, and TemporalVQA (2501.10674-imam_can_2025), TLQA (2501.07214-swetha_timelogic:_2025), REXTIME (2406.19392-chen_rextime:_2024), FrameCap (2412.02071-xue_progress-aware_2024), and VideoVista (2406.11303-li_videovista:_2024) have revealed substantial limitations in current multimodal language models, achieving accuracy significantly below human levels (15% versus 90%). These benchmarks emphasize the necessity for temporal logic annotations and highlight the difficulty models face when reasoning across video frames. In parallel, spatial datasets such as SpatialVLM (2401.12168-chen_spatialvlm:_2024), WhatsUp (2310.19785-kamath_whats_2023), DC2 (2408.15556-wang_divide_2024), and Grounded (2408.14469-chen_grounded_2024) showcase improvements through extensive spatial reasoning QA pairs and challenging spatial configurations.
Iterative and Reflective Reasoning Iterative reasoning capabilities have markedly improved through datasets like VISCO (2412.02172-wu_visco:_2024), Mulberry-260k (2412.18319-yao_mulberry:_2024), FIRE (2407.11522-li_fire:_2024), VLFeedback (2410.09421-li_vlfeedback:_2024), Silkie (2312.10665-li_silkie:_2023), TIIL (2404.18033-huang_exposing_2024), ConMe (2406.08164-huang_conme:_2024), and Reflective (2407.11422-zhang_reflective_2024), enabling models to learn from mistakes by both incorporating reflective and iterative feedback loops. For instance, VISCO demonstrated that human critiques significantly outperform model critiques (76% error correction), highlighting a persistent gap in models’ ability to independently self-correct.
Complex Multimodal Evaluation Complex multimodal datasets like EMMA (2501.05444-hao_can_2025), CoMT (2412.12932-cheng_comt:_2024), SMIR (2501.03675-li_smir:_2025), ProVision (2412.07012-zhang_provision:_2024), MAGEBench (2412.04531-zhang_magebench:_2024), MM-Vet v2 (2408.00765-yu_mm-vet_2024), JourneyBench (2409.12953-wang_journeybench:_2025), VERIFY (bi2025verify) and CompCap (2412.05243-chen_compcap:_2024) exposed limitations in current multimodal models’ integration capabilities. EMMA specifically shows that models struggle significantly with iterative multimodal interactions, particularly when tasks demand deep integration across domain-specific problems. Benchmarks like FineCops-Ref (2409.14750-liu_finecops-ref:_2025), VL-RewardBench (2411.17451-li_vlrewardbench:_2024), MM-SAP (2401.07529-wang_mm-sap:_2024), PCA-Bench (2402.15527-chen_pca-bench:_2024), AttCoSeg (2312.12423-pramanick_jack_2024), and M3CoT (2405.16473-chen_m\$^3\$cot:_2024) emphasize detailed evaluation metrics (Recall@k, AUROC). These benchmarks provide insights into models’ actual reasoning capabilities beyond traditional accuracy metrics.
Counterfactual and Logical Reasoning Datasets like CounterCurate (2402.13254-zhang_countercurate:_2024), C-VQA (2310.06627-zhang_what_2024), CausalChaos! (2404.01299-parmar_causalchaos!_2024), LogicAI (2407.04973-xiao_logicvista:_2024), and MCDGRAPH (2412.13540-zhu_benchmarking_2024) explicitly test model reasoning robustness, showing improvements through challenging counterfactual examples and structured logical questions.
Active Perception and Progressive Reasoning Active perception datasets such as ActiView (2410.04659-wang_actiview:_2024) and progressive reasoning benchmarks such as Blink (2404.12390-fu_blink:_2024), ADL-X (2406.09390-reilly_llavidal:_2024), and GUIDE (2406.18227-liang_guide:_2024) challenge models’ dynamic reasoning capabilities through active view adjustments and frame-level progression.
6 Conclusion
In this paper, we presented a comprehensive survey of existing work on reasoning in Multimodal Large Language Models (MLLMs) and Large Language Models (LLMs), with a focus on optimizing model performance and identifying the most effective reasoning trajectories. We introduced a variety of post-training and test-time computation methods, discussing their potential to enhance model capabilities on complex, real-world tasks. Our analysis highlights the critical role of reasoning in improving the visual understanding abilities of MLLMs. Furthermore, we have provided insights and outlined potential directions for future research to guide the development of more robust and efficient reasoning frameworks.
Appendix A Implications from the Learning Science Perspective
The design of CoT reasoning for LLMs aligns with several key cognitive science and learning science theories that enhance structured reasoning, problem-solving, and knowledge construction.
One key theory is Cognitive Load Theory (sweller2011cognitive; paas2004cognitive), which reduces a human’s working memory workload by breaking complex problems into manageable steps. CoT reasoning follows this approach by prompting models to generate step-by-step solutions rather than arriving at an answer in one leap. This structured decomposition mirrors how human learners handle intricate problems by offloading cognitive effort across multiple processing stages. Similarly, in the context of learning, scaffolding refers to offering temporary assistance to learners, enabling them to accomplish tasks that would be difficult or impossible for them to complete independently. This support is gradually removed as learners develop the skills and confidence required to perform the task on their own (van2010scaffolding). This supports CoT’s design by guiding LLMs through intermediate steps, much like how educators provide structured support to help students grasp new concepts before transitioning to independent problem-solving. These theories are both rooted in the Zone of Proximal Development (ZPD), a key construct in Lev Vygotsky’s theory of learning and development (vygotsky1978mind).
CoT reasoning also aligns with metacognition (lai2011metacognition) and self-explanation (bisra2018inducing; chi1994eliciting) by encouraging models to “think about their thinking”. Just as self-explanation improves human learning by prompting individuals to articulate their reasoning (chi1994eliciting; vanlehn1992model), CoT forces LLMs to justify their steps, reducing reliance on shallow heuristics. This ties into Dual-Process Theory (frankish2010dual), where CoT shifts LLMs from fast, intuitive decision-making (System 1) to deliberate, analytical reasoning (System 2), leading to more logical and consistent outputs.
Moreover, CoT fosters constructivist learning (waite2022jean; bruner2009process) by enabling LLMs to incrementally build knowledge structures through active reasoning. Instead of passively retrieving answers from training data, the model synthesizes prior knowledge with new information, improving adaptability. Additionally, analogical reasoning (gentner2017analogical) plays a role in CoT by helping LLMs map relationships between concepts, allowing them to generalize problem-solving strategies across different contexts.
By integrating these cognitive science and learning science principles, CoT reasoning enhances the interpretability, reliability, and generalization of LLM outputs, making them more aligned with human cognitive processes and educational best practices.
Inference
Adaptive Inference 2501.13926-guo_can_2025,  2501.02964-hu_socratic_2025,  2412.10471-yang_vca:_2024, 2412.02071-xue_progress-aware_2024,  2412.02172-wu_visco:_2024,  2412.03704-wang_scaling_2024, 2410.03577-zou_look_2024,  2408.17150-qu_look_2024,  2410.16198-zhang_improve_2024, 2408.02032-huo_self-introspective_2024,  2407.03008-liao_align_2024,  2406.10923-su_investigating_2024, 2406.11303-li_videovista:_2024,  2408.15556-wang_divide_2024,  2411.10440-xu_llava-cot:_2025, 2405.16071-zhao_dynrefer:_2024,  2404.06510-liao_can_2024,  2403.14743-mahmood_vurf:_2024, 2402.14767-cao_dualfocus:_2024,  2405.09711-wu_star:_2024,  2312.17240-yang_lisa++:_2024, 2311.13601-li_visual_2023,  2311.06612-pi_perceptiongpt:_2023 Reward Models 2501.13926-guo_can_2025,  2501.05444-hao_can_2025,  2411.17451-li_vlrewardbench:_2024, 2406.00645-fu_furl:_2024,  2403.12884-ke_hydra:_2024,  2410.02712-xiong_llava-critic:_2024, 2410.22315-cascante-bonilla_natural_2024,  2411.00855-cheng_vision-language_2024, 2410.16198-zhang_improve_2024, Feedback 2501.13926-guo_can_2025,  2412.02172-wu_visco:_2024,  2501.02964-hu_socratic_2025, 2503.07365-meng_mm-eureka:_2025,  2407.11522-li_fire:_2024,  2410.02712-xiong_llava-critic:_2024, 2411.00855-cheng_vision-language_2024,  2410.16198-zhang_improve_2024,  2411.12591-zheng_thinking_2024, 2404.06510-liao_can_2024,  2402.01345-han_skip_2024,  2405.19209-wang_videotree:_2024, 2403.16999-shao_visual_2024,  2404.04007-liang_neural-symbolic_2024, 2311.07362-lee_volcano:_2024,  2312.03052-hu_visual_2024, Multimodal Decomposition 2503.06252-xiang_can_2025,  2501.13536-liang_reasvqa:_2025,  2501.05069-liu_commonsense_2025, 2412.08564-shlapentokh-rothman_can_2024,  2409.19339-zhang_visual_2024, 2408.14469-chen_grounded_2024,  2409.08202-hsu_what_2024,  2406.10923-su_investigating_2024, 2406.09390-reilly_llavidal:_2024,  2408.05019-wang_instruction_2024,  2406.16620-zhang_omagent:_2024, 2404.06511-min_morevqa:_2024,  2404.01299-parmar_causalchaos!_2024, 2403.09394-wang_git:_2024,  2404.01911-dzabraev_vlrm:_2024,  2405.09711-wu_star:_2024, 2312.14135-wu_v*:_2023,  2309.15112-zhang_internlm-xcomposer:_2023, Efficiency and Scalability 2502.00382-goyal_masked_2025,  2501.07542-li_imagine_2025,  2412.08635-sun_multimodal_2024, 2412.18072-fan_mmfactory:_2024,  2501.02669-park_generalizing_2025,  2408.15556-wang_divide_2024, 2409.12953-wang_journeybench:_2025,  2408.02032-huo_self-introspective_2024, 2407.04681-lin_rethinking_2024,  2410.03321-ni_visual-o1:_2024,  2405.19209-wang_videotree:_2024, 2406.08164-huang_conme:_2024,  2403.07487-zhang_motion_2024,  2403.09333-zhan_griffon_2024, 2403.19322-chen_plug-and-play_2024,  2403.09394-wang_git:_2024,  2312.11420-zhao_tuning_2023, 2311.06612-pi_perceptiongpt:_2023,  2401.12168-chen_spatialvlm:_2024 Search Strategies 2412.18319-yao_mulberry:_2024,  2412.10471-yang_vca:_2024,  2412.03704-wang_scaling_2024, 2501.10674-imam_can_2025,  2411.10440-xu_llava-cot:_2025,  2408.17150-qu_look_2024, 2408.15556-wang_divide_2024,  2406.11303-li_videovista:_2024,  2406.09175-kazemi_remi:_2024, 2404.15190-shin_socratic_2024,  2405.03272-zhang_worldqa:_2024, 2405.16071-zhao_dynrefer:_2024,  2406.00645-fu_furl:_2024,  2403.14743-mahmood_vurf:_2024, 2312.14135-wu_v*:_2023,  2306.08129-hu_avis:_2023,
Figure 4: Comprehensive Overview of Methods and Frameworks focus on test-time compute
Post
Training
Policy Optimization 2406.00645-fu_furl:_2024,  2402.07384-zhang_exploring_2024,  2312.03052-hu_visual_2024, 2310.07932-tian_what_2024,  2309.10790-kim_guide_2023,  2402.06118-yan_vigor:_2024, 2408.10433-ouali_clip-dpo:_2024,  2410.02712-xiong_llava-critic:_2024,  2410.04659-wang_actiview:_2024, 2501.18880-waite_rls3:_2025,  2503.08525-wei_gtr:_2025,  2503.05379-zhao_r1-omni:_2025, 2411.10440-xu_llava-cot:_2025,  2410.17885-deng_r-cot:_2024, Model Architecture 2406.09390-reilly_llavidal:_2024,  2404.13847-ma_eventlens:_2024,  2404.01258-zhang_direct_2024, 2403.09333-zhan_griffon_2024,  2401.17981-jiao_training-free_2024, 2401.03105-he_incorporating_2024,  2312.08870-ma_vista-llama:_2023,  2312.03631-ben-kish_mitigating_2024, 2312.03052-hu_visual_2024,  2311.03079-wang_cogvlm:_2024,  2410.09575-wang_reconstructive_2024, 2410.08209-cao_emerging_2024,  2408.14469-chen_grounded_2024,  2408.13890-huang_making_2024, Reward Alignment 2501.05444-hao_can_2025,  2411.17451-li_vlrewardbench:_2024,  2410.02712-xiong_llava-critic:_2024, 2403.12884-ke_hydra:_2024,  2406.00645-fu_furl:_2024,  2404.01911-dzabraev_vlrm:_2024, 2312.03052-hu_visual_2024, Iterative Refinement 2503.07365-meng_mm-eureka:_2025,  2411.12591-zheng_thinking_2024,  2404.06510-liao_can_2024, 2412.02172-wu_visco:_2024,  2403.16999-shao_visual_2024, 2402.01345-han_skip_2024,  2312.03052-hu_visual_2024,  2311.07362-lee_volcano:_2024, Spatial Temporal Modeling 2405.19209-wang_videotree:_2024,  2403.07487-zhang_motion_2024,  2401.12168-chen_spatialvlm:_2024, 2312.14135-wu_v*:_2023, Dataset Curation 2412.18072-fan_mmfactory:_2024,  2409.12953-wang_journeybench:_2025, 2404.16222-nagarajan_step_2024,  2405.03272-zhang_worldqa:_2024, 2407.04681-lin_rethinking_2024,  2312.11420-zhao_tuning_2023,  2306.08129-hu_avis:_2023,
Figure 5: Comprehensive Overview of Methods and Frameworks focus on post-training improvement